If (a-b) is 6 more than (c+d) and (a+b) is 3 less than (c-d), then the value of (a-c) is
Answer by tommyt3rd(5050)

If (a-b) = 6 + (c+d)
and
(a+b) = (c-d) - 3

(a-b)-(c+d) = 6
(a+b)-(c-d)= - 3
2a-2c=3
a-c=3/2